What is the equation of the line that passes through (4, -1) and (-2, 3)?
Can you help me answer this? I keep getting y = -4/6x + ? and it has to be written in general form.

Answer by Alan3354(69443)

What is the equation of the line that passes through (4, -1) and (-2, 3)?
Can you help me answer this? I keep getting y = -4/6x + ? and it has to be written in general form.
-------------------------
Find the slope, m
m = diffy/diffx = (3 - (-1))/(-2 -4) = 4/-6 = -2/3
----------
Use y = mx + b and either point to find b
-1 = (-2/3)*4 + b
b = -1 + 8/3 = 5/3
-------------
Then it's y = mx + b
--------------------
The eqn is
y = (-2/3)x + 5/3
Or 2x + 3y = 5